y"+3y'+2y=e^(-t) , y(0)=0, y'(0)=0


1
Expert's answer
2022-04-07T06:42:52-0400

"y''+3y+2y=e^{-t},y\\left( 0 \\right) =0,y'\\left( 0 \\right) =0\\\\\\lambda ^2+3\\lambda +2=0\\Rightarrow \\lambda \\in \\left\\{ -1,-2 \\right\\} \\\\y=C_1e^{-t}+C_2e^{-2t}-the\\,\\,solution\\,\\,of\\,\\,homogeneous\\,\\,equation\\\\The\\,\\,partial\\,\\,solution\\,\\,of\\,\\,the\\,\\,form\\,\\,Ate^{-t},\\sin ce\\,\\,-1 is\\,\\,a\\,\\,characteristic\\,\\,number\\,\\,of\\,\\,multiplicity\\,\\,1\\\\y'=\\left( -At+A \\right) e^{-t}\\\\y''=\\left( At-2A \\right) e^{-t}\\\\\\left( At-2A \\right) e^{-t}+3\\left( -At+A \\right) e^{-t}+2Ate^{-t}=e^{-t}\\\\Ae^{-t}=e^{-t}\\Rightarrow A=1\\\\Thus\\,\\,\\\\y=te^{-t}+C_1e^{-t}+C_2e^{-2t}\\\\y'=e^{-t}-te^{-t}-C_1e^{-t}-2C_2e^{-2t}\\\\\\left\\{ \\begin{array}{c}\ty\\left( 0 \\right) =0\\\\\ty'\\left( 0 \\right) =0\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tC_1+C_2=0\\\\\t1-C_1-2C_2=0\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tC_1=-1\\\\\tC_2=1\\\\\\end{array} \\right. \\\\y=te^{-t}-e^{-t}+e^{-2t}"